OpenStudy (anonymous):

A line has a slope of 1/2 and contains the points (3, y) and (5, 7). The value of y is A.8 B.6 C.3

OpenStudy (anonymous):

First off, the slope is rise/run, meaning, y/x

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

5/4=>-4/5 use point slope formula y-y1=m(x-x1) y--2=(-4/5)(x-4) y+2=-4/5x+16/5 y=-4/5x+6/5 <=slope form standard form: 4x+5y=6

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

wc

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

i thought it was c

OpenStudy (anonymous):

I'm trying to bump my question, why is it not working xD

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

If it was c that would make the slope 2/1 not 1/2

OpenStudy (anonymous):

So it is b

OpenStudy (anonymous):

oh wow i mix up the variables.